OpenStudy (fools101):

Logarithmc Equations logv2 (9x + 5) - logv2 (x^2 - 1) = 2

OpenStudy (fools101):

Please help!

OpenStudy (anonymous):

what is v doing above?

OpenStudy (fools101):

 IT that sorry!

OpenStudy (anonymous):

3,-3/4

OpenStudy (anonymous):

by law of logs logv2 (9x + 5) - logv2 (x^2 - 1) = log2 [(9x+5)/(x^2-1) = 2 (9x+5)/(x^2-1) = 2^2 = 4 9x+ 5 = 4x^2 - 4 4x^2 - 9x - 9 = 0 (4x + 3 )(x - 3) = 0 x = -3/4 , 3

OpenStudy (anonymous):

the only valid root is 3 as -3/4 makes log (9x+5) = log of negative which does not exist

OpenStudy (fools101):

Thank u!